Suppose 1 year and 2 year bond yields currently are 2% and 3%. According to expectations hypothesis, whats the expected yielPlease explain how it is 4%
Answer #1

Based on Expectations hypothesis,

(1 + 2 Yr Yield)2 = (1 + 1 Yr Yield) * (1 + 1 Yr Yield 1 Yr from now)

(1 + 3%)2 = (1 + 2%) * (1 + 1 Yr Yield 1 Yr from now)

1.0609 = 1.02 * (1 + 1 Yr Yield 1 Yr from now)

(1 + 1 Yr Yield 1 Yr from now) = 1.040098

1 Yr Yield 1 Yr from now = 0.040098

1 Yr Yield 1 Yr from = 4.00%